OpenStudy (anonymous):

can someone show me how to do this? : 3. Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.

OpenStudy (31356):

w(w^2-w-2)-40=0 Solving that is faster if you know synthetic division. You would use the idea of the Rational Roots Theorem to find values for w to satisfy the cubic equation. I'd suggest first working with +/-4, +/-5, +/-8, +/-10. When you get just one first root, the next two will be easy. Some further work: Good News! I tried some long divisions and found +5 or -5 are not roots, but that +4 is a root. One of the binomials is (w-4). The quotient from this was w%5E2%2B3w%2B10. So this means your polynomial equation is %28w-4%29%28w%5E2%2B3w%2B10%29=0, so the quadratic part should be much easier to manage. NOTE: The discriminant for that quadratic is -31, so the solution will contain an imaginary part. The only reasonable value for w is 4.

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

what is with the %28w-4%29%28w%5e2%2b3 ect

OpenStudy (31356):

That is the quotient with the percents and w means width.

OpenStudy (anonymous):

i find that very hard to read sorry, i know about w+1 and w-2 as far as i know i would just have to find the first root by solveing (w^2 +w)(w-2) = 40

OpenStudy (31356):

Sorry about that, that's all I can explain or help you with.

OpenStudy (anonymous):

kk thx ill ask on Yahoo!

OpenStudy (31356):

l * w * h = 40 l = w + 1 h = w - 2 Substitute 2 and 3 into 1 (w + 1) * w * (w - 2) = 40 w^3 - w^2 -2w = 40 w = 4 l = 5 h = 2

OpenStudy (31356):

Here's one I found on Yahoo!

OpenStudy (31356):

Medal?

OpenStudy (31356):

I guess that is answer. LOL

OpenStudy (31356):

I guess that is the answer. LOL

OpenStudy (31356):

;D